OpenStudy (anonymous):

Use the principle of square roots. x^2+13/3x+169/36=13/36

OpenStudy (anonymous):

x^2 + 13/3 x + 169/36 = 13/36\[\frac{1}{36} (13+6 x)^2=13/36 \]\[(13+6 x)^2=13 \]\[\left\{x\to \frac{1}{6} \left(-13-\sqrt{13}\right),x\to \frac{1}{6} \left(-13+\sqrt{13}\right)\right\} \]